### v1

# Revealing Trends in Datasets from the 2022 ACL and EMNLP Conferences

Abstract
Natural language processing (NLP) has grown significantly since the advent of the Transformer architecture. Transformers have given birth to pre-trained large language models (PLMs). There has been tremendous improvement in the performance of NLP systems across several tasks. NLP systems are on par or, in some cases, better than humans at accomplishing specific tasks. However, it remains the norm that better quality datasets at the time of pretraining enable PLMs to achieve better performance, regardless of the task. The need to have quality datasets has prompted NLP researchers to continue creating new datasets to satisfy particular needs. For example, the two top NLP conferences, ACL and EMNLP, accepted ninety-two papers in 2022, introducing new datasets. This work aims to uncover the trends and insights mined within these datasets. Moreover, we provide valuable suggestions to researchers interested in curating datasets in the future.
1 Introduction
Natural language processing (NLP) has grown significantly since the advent of the Transformer architecture. Transformers have given birth to a new pre-trained of large language models (PLMs).
There has been tremendous improvement in the performance of NLP systems across several tasks. NLP systems are on par or, in some cases, better than humans at accomplishing specific tasks.
However, it remains the norm that better quality datasets at the time of training enable PLMs to perform better, regardless of the task. The need to have quality datasets has prompted NLP researchers to continue creating new datasets to satisfy particular needs. For example, the two top NLP conferences, ACL and EMNLP, accepted ninety-two papers in 2022, introducing new datasets. This work aims to uncover the trends and insights mined within these datasets. Figure 1 shows the main NLP tasks for which authors created new datasets, which were published at both EMNLP and ACL in the year 2022. The red box in Figure 1 indicates the tasks in both ACL and EMNLP datasets.
We draw inspiration from Gollapalli and Li (2015), who previously studied the significant themes across NLP conferences from 1996 to 2014. The more recent work to examine the disparate sources utilized to construct new datasets is the study by Yu et al. (2022). We aim to reveal insights such as significant author affiliations in creating datasets for training NLP models.
Lastly, we provide helpful suggestions to researchers interested in curating datasets in the future.
Figure 1: An overview of major NLP tasks covered in datasets published in ACL and EMNLP in 2022.
Figure 2: Datasets, major NLP tasks, and author-affiliation for datasets published at ACL and EMNLP in 2022.
2 Scope of this Study
This paper’s analysis and findings are based only on documents accepted at the two top-tier NLP conferences, ACL2022 and EMNLP2022. These conferences are a good measure of contemporary NLP needs; thus, the work published in these two conferences best reflects the trajectory in which NLP research is headed.  [^1] The LREC conference is a significant collection of datasets, but we aim at conferences considered top tier in NLP.
Article Retrieval
To get the relevant papers during this study, we used the keywords dataset and corpus. Initially, we obtained 130 papers. We then verified that each article introduced a new dataset or corpus. After this inspection, only 92 papers remained. The resulting papers were scrutinized to extract the following attributes: name of dataset, dataset size, NLP topic, primary baselines, multilingualism, multimodality, and author-affiliation. These attributes were chosen to extract as much relevant information as possible, the kind of information that would permit us to understand the recent trend in curating new NLP datasets.
3 NLP Tasks Coverage in Datasets
First, we set out to categorize the significant tasks represented in the datasets to provide a candid analysis of the datasets. Figure 1 shows the significant topics that are covered by the datasets published at ACL—in the green box; and topics covered by datasets published at EMNLP—in the blue box. Topics common to ACL and EMNLP are shown in the middle—in the red box.
In addition, from both Table 1 and Table 2, it is evident that text generation, text summarization, text or token classification, information extraction, natural language understanding, question answering are the focal point for dataset creation for NLP researchers.
Major topics in ACL 2022
Dialogue Reasoning, Document Retrieval, Reading Comprehension, Question Answering, Causal Reasoning, Bias Detection, Natural Language Generation, Argument Mining, Legal NLU, Multimodal chat translation, Multimodal dialogue sentiment analysis, Text Generation, Abstractive Summarization, Genomics Data Summary Generation, Hate Speech Detection, Automated Fact Checking, Explainable Feedback, Multilingual Relation Extraction, Automatic Speech Recognition, Sign Language Recognition, Chinese Synesthesia Detection, Visual Question Answering, Cultural Background Prediction, Dialogue, Legal Event Detection, Controllable Text Generation, Dialogue Safety, Multilingual Document Understanding, Semantic Entity Recognition
Table 1: The NLP topics in datasets at ACL 2022.
Major topics in EMNLP 2022
Multilingual image captioning, Coreference Resolution, Event coreference, Temporal, Causal, and Subevent Relation Extraction, Multi-Domain Language Modeling, Model bias, Gender accuracy in translation, Open domain chatbots, Drug safety, Named Entity Recognition, Event extraction, Figurative Language Identification, Paraphrase identification and generation, Extractive Summarization, Dialogue Summarization, Question Matching, Hate Speech in Dialogue, Language Modeling, Image captioning, Visual storytelling, Speech Recognition, Reading Comprehension, Reasoning, Data Augmentation, Relational Extraction, Sequence Labeling, Offensive Language detection, Dialog Generation, Patent Classification, Zero-shot Learning, Multimodal dialogue summarization, response generation, Topic Modeling, Commonsense knowledge, Movie Understanding and Narrating, Visual Entity Linking, Question Answering, Multilingual Information Extraction, Natural Language Inference, Text-to-SQL, Politeness Classification, Graphs Extraction, Source code generation, Post-OCR text correction, Hate Speech Detection, Event Extraction
Table 2: The NLP topics in datasets at EMNLP 2022.
Table 1 indicates the list of the significant NLP topics tasks represented in the papers that introduced datasets and were published at ACL in 2022. Similarly, the NLP topics covered in the datasets published at EMNLP 2022 are shown in Table 2.
4 Dataset size
The number of sentences in datasets published at ACL and EMNLP is essential to understanding the number of sentences required to build an NLP system with satisfactory performance.
Data size (# instances) Dataset Count
$\leq$ 500 3
500 - 1,000 2
1,000 - 1,500 3
1,500 - 2,000 1
2,000 - 3,000 2
3,000 - 4,000 5
4,000 - 5,000 4
5,000 - 10,000 8
10,000 - 50,000 24
50,000 - 100,000 5
100,000 - 500,000 13
500,000 - 1,000,000 0
$\geq$ 1,000,000 3
Table 3: Dataset size at ACL and EMNLP in 2022.
We have summarized the size of datasets in Table 3 from which it is evident that most datasets have data samples in the range of 10,000 to 50,000.
5 Collaboration in Dataset Construction
It is important to list all the affiliations of authors at both conferences to understand the level of collaboration, especially between academic and industry researchers.
Author Affiliation at ACL 2022
University of Michigan, Singapore University of Technology and Design, Maastricht University, Carnegie Mellon University, Google Research, Harbin Institute of Technology, University of California Irvine, IBM Research, Tencent, Rensselaer Polytechnic Institute, University of Washington, University of Notre Dame, Syracuse University, Hong Kong Univ. of Sci and Tech, Georgia Institute of Technology, University of California Los Angeles, Columbia University, IBM Research Ireland, Université Paris-Saclay, ObTIC, Sorbonne Université, Université de Lorraine, Shanghai Jiao Tong University, Microsoft Research Asia, Xi’an Jiaotong University, DAMO Academy Alibaba Group, The Chinese University of Hong Kong, National University of Singapore, Tsinghua University, Cloud BU Huawei Technologies, ETH Zurich, Noah’s Ark Lab Huawei Technologies, Nanyang Technological University, University of Copenhagen, Universitat Hamburg, Bucerius Law School, Stanford Law School, Athens University of Economics and Business, University of Sheffield, Chicago Kent College of Law, Beijing Jiaotong University, University of Tokyo, RIKEN AIP, Yahoo Japan, Toyota Technological Institute at Chicago, Duke University, University of Science and Technology of China, MIT, Microsoft Research, Allen Institute for AI, Microsoft, University of Waterloo, Fudan University, Shopee-Singapore, Vipshop (China) Co.Ltd, Fudan-Aishu Cognitive Intelligence Joint Research Center, Technical University of Darmstadt, IIT Delhi, Babylon Health, University of Aberdeen, University of Manchester, Soochow University, Chinese Academy of Social Sciences, University of Essex, Baidu Inc, Dartmouth College, California Institute of Technology, UC Santa Barbara, Vector Institute, Beijing Academy of Artificial Intelligence, Beijing Powerlaw Intelligent Technology Co, LMU Munich, Microsoft Azure AI, WeChat AI Tencent, Tencent AI Lab
Table 4: At ACL 2022, there were 66 unique institutions to which dataset authors are affiliated.
Author Affiliation at EMNLP 2022
Google Research, Bar Ilan University, Intel Labs Israel, Salesforce AI Research, Shanghai Mental Health Center, University of Cambridge, AWS AI Labs, Yonsei University, Seoul National University, University of Warwick, Northeastern University, AstraZeneca, King’s College London, The Alan Turing Institute, Institut Polytechnique de Paris, Télécom-Paris, École Polytechnique, Korea University, Adobe Research, AIGEN Sciences, Hong Kong University of Science and Technology, University of Science and Technology of China, Microsoft Azure Cognitive Services Research, Heidelberg University, SRH Hochschule Heidelberg, Beijing University of Posts and Telecommunications, Soochow University, Fudan University, Baidu Inc, University of Trento, Free University of Bozen-Bolzano, Fondazione Bruno Kessler, University of Dayton Research Institute, SIL International, Masakhane, Carnegie Mellon University, Microsoft Semantic Machines, University of Utah, Meta AI, University of Oregon, Adobe Research, UC San Diego, UMass Amherst, DAMO Academy Alibaba Group, Singapore University of Technology and Design, Hong Kong Polytechnic University, KAIST, SoftlyAI Research, IIT Kharagpur, Goldman Sachs Data Science and ML Group, Leibniz University of Hannover, IIT Patna, Bosch Center for Artificial Intelligence, Heidelberg University, Shanghai AI Laboratory, Huawei Noah’s Ark Lab, University of Washington, The University of Hong Kong, JD AI Research, Umeå University, Adlede AB, Zhejiang University, AZFT Joint Lab for Knowledge Engine, UNC Chapel Hill, Allen Institute for AI, Tencent AI Lab, Renmin University of China, Hong Kong Polytechnic University, Chinese Academy of Sciences, University of Chinese Academy of Sciences, Islamic University of Technology, Chinese University of Hong Kong, Tsinghua University, IT University of Copenhagen, LMU Munich, Munich Center for ML, Shanghai AI Lab, University of Melbourne, RMIT University, MBZUAI, Data Friendly Space, ToggleCorp Solutions, Johannes Kepler University Linz, Stony Brook University, University of Arizona, Stanford University, VinAI Research, Hanoi University of Science and Technology, University of Oxford, University of Texas at Austin, UC Santa Barbara, Southeast University, Nanjing Origin Information Technology Company, IIT Bombay, Uniphore, HuggingFace, Yale University, University of Southern California, Kakao Corp, Soongsil University, Seoul National University, Virginia Tech, Jamia Millia, University of Sydney, University of New South Wales, University of Notre Dame
Table 5: At EMNLP 2022, there were 103 unique institutions to which dataset authors are affiliated.
Table 4 and Table 5 indicate all the institutions, academic or industry, to which the authors are affiliated.
The big winners in academia include Tsinghua University, the University of Washington, the Singapore University of Technology and Design, the National University of Singapore, Nanyang Technological University, the Hong Kong University of Science and Technology, and Stanford.
The industry’s big winners include Microsoft Research, Adobe Research, Google Research, Huawei Noah’s Ark Lab, Alibaba DAMO Academy, and Tencent AI Lab.
Collaboration
There are 26 papers in which all the authors belong to one institution, either academic or industry. Moreover, the number of documents in which all the authors are from one academic institution is 20.
Most papers consist of authors hailing from both academia and industry. This points to the advantages enjoyed when industry and academia researchers collaborate. It re-emphasizes long-held beliefs that the industrial labs provide the practical use cases, large-scale data, computing resources, and funds to foot the costs necessary to build a new dataset. In contrast, academia provides theoretical insights, novel methodologies, and expertise in meticulous experimental design. Out of the forty-three papers in which the industry and academic researchers collaborated, the number of academic last authors is 20 while industry last authors is 23. There are also few academia-academia collaborations.
6 Major Baselines
We discuss contemporary baselines for major tasks. Full details in Table 6.
NLP Task Baselines
Question Answering BERT, DistilBERT, BART
Abstractive Summarization Transformer, BART, BM25, T5, PEGASUS, Longformer
Extractive Summarization BERT, RoBERTa, Longformer
Event Extraction Pipeline approach: OneIE, FourIE; and for Joint approach: BERT, RoBERTa
Relation Extraction BERT, RoBERTa, SciBERT, PubmedBERT, BioBERT, BlueBERT
Coreference Resolution, Event Coreference RoBERTa, Deep Passage Retrieval
Text Generation, NLG LexRank, TextRank, MMR, LSTM-CRF, BART
Hate Speech Detection HateBert, ToxDectRoBERTa and LSTM, BERT, RoBERTa, DeBERTa
Natural Language Understanding TFIDF-SVM, BERT, RoBERTa, DeBERTa, Longformer, BigBird
Named Entity Recognition BERT, RoBERTa
Machine translation mBERT, mBART, mT5, OPUS-MT
Image captioning ViT, mT5, CIDEr, InceptionV3
Table 6: The major baselines from papers in our study.
7 The Rise of Multimodality
The growing demand for NLP systems that accept at least two modalities as input, for example, text and image, leads to a further increase in demand for multimodal datasets.
Such multimodal datasets make it possible to meet the needs for both textual and visual embeddings when training visual-language systems for tasks such as multimodal dialogue summarization, multimodal response generation, visual question answering, multimodal dialogue sentiment analysis, visual storytelling, visual entity linking, multimodal information extraction, among other things.
Papers which introduced multimodal datasets include: MSCTD (Liang et al., 2022), WikiDiverse (Wang et al., 2022f), DuReader_vis (Qi et al., 2022), XFUND (Xu et al., 2022a), Crossmodal-3600 (Thapliyal et al., 2022), FigMemes (Liu et al., 2022b)(Leong et al., 2022), JDDC 2.1 (Zhao et al., 2022c),  Sun et al. (2022c), and (Thapa et al., 2022).
MSCTD (Liang et al., 2022) consists of 142,871 English-Chinese utterance pairs and supports the development of multimodal chat translation and dialogue sentiment analysis. WikiDiverse (Wang et al., 2022f) contains 8K image captions and makes evaluating multimodal entity linking (MEL) systems possible. DuReader_vis (Qi et al., 2022) includes 15K QA pairs in Chinese and 158K document images, making evaluating visual question answering (VQA) systems possible. XFUND (Xu et al., 2022a) introduced 1,393 forms and supports multilingual document understanding, semantic entity recognition, and relation extraction for seven languages, that is, Chinese, Japanese, Spanish, French, Italian, German, and Portuguese. Crossmodal-3600 (Thapliyal et al., 2022) comprises 3,600 images and 261,373 image captions necessary to support the development of multilingual image captioning systems. FigMemes (Liu et al., 2022b), which contains 5,141 labels, makes it possible to identify figurative language in politically-opinionated memes. BLOOM (Leong et al., 2022) is a massive dataset that covers 363 languages and supports the development of image captioning, visual storytelling, speech synthesis and recognition, and language modeling. On the other hand, JDDC 2.1 (Zhao et al., 2022c) enables the development of query writing, response generation, discourse parsing, and summarization tasks in Chinese. MovieUN (Zhang et al., 2022) enables movie understanding in Chinese. To allow the development of visual entity linking systems, Sun et al. (2022c) developed a new dataset. Thapa et al. (2022)’s dataset enables hate speech detection in natural language.
8 Quantifying the Language Gap: Monolingual vs. Multilingual Data
Unsurprisingly, most datasets published at ACL and EMNLP conferences contain English text. However, there is a notable number of non-English datasets, too. To study the multilingual aspect of these datasets, we categorized the datasets into (i) Non-English and Monolingual, which includes only datasets curated with data in one language and that language is not English; and (ii) Multilingual in which the data is curated in more than one language.
We have provided further details for all the datasets in the Appendix B.
BanglaFrenchKoreanPatois*Chinese001010101011112222222211118888Language in the dataset.Count.Monolingual data
Figure 3: The number of non-English monolingual datasets. Patois* refers to Jamaican Patois.
Non-English and Monolingual
This category includes datasets containing only one language and that language is not English. Refer to Table 7.
The five languages represented in this category are: French—French CrowS-Pairs (Névéol et al., 2022) and Louis and Spanakis (2022); Bangla—BanglaRQA (Ekram et al., 2022); Korean—KOLD (Jeong et al., 2022) and APEACH (Yang et al., 2022a); Jamaican Patois—JamPatoisNLI (Armstrong et al., 2022); Chinese— Jiang et al. (2022), DuReader_vis (Qi et al., 2022), LEVEN (Yao et al., 2022b), D4 (Yao et al., 2022a), Title2Event (Deng et al., 2022), DuQM (Zhu et al., 2022),  (Zhao et al., 2022b), JDDC 2.1 (Zhao et al., 2022c), MovieUN (Zhang et al., 2022) and  (Zhou et al., 2022).
Moreover, Maheshwari et al. (2022) introduced a benchmark dataset for Sanskrit.
Monolingual Dataset Language
French CrowS-Pairs (Névéol et al., 2022) and  Louis and Spanakis (2022) French
BanglaRQA (Ekram et al., 2022) Bangla
 Maheshwari et al. (2022) Sanskrit
KOLD (Jeong et al., 2022) and APEACH (Yang et al., 2022a) Korean
JamPatoisNLI (Armstrong et al., 2022) Jamaican Patois
Jiang et al. (2022), DuReader_vis (Qi et al., 2022), D4 (Yao et al., 2022a), Title2Event (Deng et al., 2022), DuQM Zhu et al. (2022)(Zhao et al., 2022b), JDDC 2.1 (Zhao et al., 2022c), MovieUN (Zhang et al., 2022),  (Zhou et al., 2022) Chinese
Table 7: The Non-English but monolingual datasets.
Data Source Tests Legal Govt Book Crowdsourcing Online Movie,TV Prompting News Med. OldData
Count 1 3 2 4 15 31 2 2 7 3 6
Table 8: Sources of text, image, and video used to construct datasets. Test* refers to language test materials. Legal* refers to legal documents e.g., law articles and patents. Govt means government websites. Online* includes forums, GitHub, e-commerce platforms, search engines, and social media. Med* refers to medical reports and biomedical literature. oldData* refers to existing datasets modified to create new datasets.
Multilingual
These datasets contain multiple languages (see Table 9). MSCTD (Liang et al., 2022) in English and Chinese; Filighera et al. (2022) in English and German; DiS-ReX (Bhartiya et al., 2022) in English, French, Spanish and German; Maronikolakis et al. (2022) in Brazilian Portuguese, German, Hindi, Swahili and English; XFUND (Xu et al., 2022a) in 7 languages Chinese, Japanese, Spanish, French, Italian, German and Portuguese; Crossmodal-3600 (Thapliyal et al., 2022) in 36 languages; MT-GenEval Currey et al. (2022) in Arabic, French, German, Hindi, Italian, Portuguese, Russian and Spanish; IndicNLG Benchmark Kumar et al. (2022) in 11 Indic languages; EUR-Lex-Sum (Aumiller et al., 2022) in 24 official European languages; ClidSum (Wang et al., 2022b) in English, German and Chinese; BLOOM (Leong et al., 2022) in 363 languages; MEE (Pouran Ben Veyseh et al., 2022) in 8 languages English, Spanish, Portuguese, Polish, Turkish, Hindi, Korean, and Japanese; HumSet (Fekih et al., 2022) in English, Spanish, French;  Lai et al. (2022) in 5 languages English, Danish, Spanish, Turkish, and Urdu; TyDiP (Srinivasan and Choi, 2022) in 9 languages Hindi, Korean, Spanish, Tamil, French, Vietnamese, Russian, Afrikaans, Hungarian; and ClassBases (Wiriyathammabhum, 2022) in English, Spanish, Portuguese.
Multilingual Dataset # Lang. Language
MSCTD (Liang et al., 2022) 2 English, Chinese
Filighera et al. (2022) 2 English, German
ClidSum (Wang et al., 2022b) 3 English, German, Chinese
ClassBases (Wiriyathammabhum, 2022) 3 English, Spanish, Portuguese
HumSet (Fekih et al., 2022) 3 English, Spanish, French
DiS-ReX (Bhartiya et al., 2022) 4 English, French, Spanish, German
Maronikolakis et al. (2022) 5 Brazilian Portuguese, German, Hindi, Swahili, English
Lai et al. (2022) 5 English, Danish, Spanish, Turkish, Urdu
XFUND (Xu et al., 2022a) 7 Chinese, Japanese, Spanish, French, Italian, German, Portuguese
MEE (Pouran Ben Veyseh et al., 2022) 8 English, Spanish, Portuguese, Polish, Turkish, Hindi, Korean, Japanese
MT-GenEval (Currey et al., 2022) 8 Arabic, French, German, Hindi, Italian, Portuguese, Russian, Spanish
TyDiP (Srinivasan and Choi, 2022) 9 Hindi, Korean, Spanish, Tamil, French, Vietnamese, Russian, Afrikaans, Hungarian
IndicNLG Benchmark (Kumar et al., 2022) 11 Assamese, Bengali, Gujarati, Hindi, Marathi, Odiya, Punjabi, Kannada, Malayalam, Tamil, Telugu
EUR-Lex-Sum (Aumiller et al., 2022) 24 Bulgarian, Croatian, Czech, Danish, Dutch, English, Estonian, Finnish, French, German, Greek, Hungarian, Irish, Italian, Latvian, Lithuanian, Maltese, Polish, Portuguese, Romanian, Slovak, Slovenian, Spanish, Swedish
Crossmodal-3600 (Thapliyal et al., 2022) 36 Arabic, Bengali, Chinese-Simplified, Croatian, Cusco Quechua, Czech, Danish, Dutch, English, Filipino, Finnish, French, German, Greek, Hebrew, Hindi, Hungarian, Indonesian, Italian, Japanese, Korean, Maori, Norwegian, Persian, Polish, Portuguese, Romanian, Russian, Spanish, Swahili, Swedish, Telugu, Thai, Turkish, Ukrainian, Vietnamese.
BLOOM Leong et al. (2022) 363 Full list of 363 languages in Appendix C.
Table 9: Multilingual datasets in order of increasing number of languages.
9 Source of Data, Dataset Creation
A thorough analysis of the sources of texts and images (see Table 8) used to construct the datasets reveals the following. 1) Diverse online resources play a critical role in availing the texts, images, and videos leveraged to construct datasets. The online sources vary from GitHub repositories, online forums, search engines and e-commerce, e.g., Baidu and JD.com; social media platforms, e.g., Reddit and Twitter; digital book collections; video and TV streaming platforms, e.g., YouTube and Xigua Video, and NGO and government websites. 2) Wikipedia and news articles remain a profound source of texts needed to construct new corpora. 3) The materials used to construct datasets include English language exams, patents, news articles, textbooks, medical reports, image collections, and movie clips. 4) Crowdsourcing, which involves recruiting human annotators, remains a relevant method to collect data. 5) While most datasets introduce entirely new data, some authors re-purpose existing datasets, creating new corpora that suit new tasks or analyses. 6) We note that prompting large language models (LLMs) to generate training examples has emerged as a new method to overcome the time and budget constraints that come with recruiting human annotators.
10 Conclusion
Our study focuses exclusively on datasets and NLP methodologies presented at the ACL and EMNLP conferences in 2022, selected for their outstanding quality and influential insights in natural language processing. These conferences are recognized for setting trends and directions in NLP research, ensuring that our analysis reflects the forefront of the discipline. Despite the narrow scope, we anticipate our findings will offer substantial value by highlighting innovative datasets introduced at these venues, serving as a vital resource and guide for researchers aiming to develop new datasets in NLP.
Acknowledgements
We thank Iqra Ali from the NAIST-NLP group for helping with gathering information about the data sources.
Appendix A Dataset Size
0.5K1K1.5K2K3K4K5K10K50K100K500K1M>1M0088881616161624242424323232323333222233331111222255554444888824242424555513131313003333Dataset SizeCountDataset size
Figure 4: The varying sizes of datasets. Most datasets tend to have between 10K—50K samples.
Appendix B Dataset details
Details for all the datasets surveyed. Due to a space limitation, we split the table into seven parts. Therefore, the Tables 10, 11, 12, 13, 14, 15, 16 are a presented in a continuous manner.
Paper/ Dataset Size (#sentences) NLP Topic Baselines Multi-lingual? Multi-modal?
Track: ACL 2022
 Ghosal et al. (2022) CICERO: A Dataset for Contextualized Commonsense Inference in Dialogues 53,105 inst. dialogue reasoning GLUCOSE-T5, RoBERTa-large, ELECTRA-large, T5-large, Unified QA Large No No
 Louis and Spanakis (2022) A Statutory Article Retrieval Dataset in French 1,100 Statutory article retrieval TF-IDF, BM25, word2vec, fastText, RoBERTa, CamemBERT No (French) No
 Sun et al. (2022a) ConditionalQA: A Complex Reading Comprehension Dataset with Conditional Answers 3,427 questions Reading Comprehension, QA ETC, DocHopper, FiD, Human No No
 Du et al. (2022) e-CARE: a New Dataset for Exploring Explainable Causal Reasoning 21,000 Causal Reasoning BERT, RoBERTa, XLNet, ALBERT, GPT2, BART No No
 Xu et al. (2022b) Fantastic Questions and Where to Find Them: FairytaleQA – An Authentic Dataset for Narrative Comprehension 10,580 questions Question answering BERT, DistilBERT, BART No No
 Névéol et al. (2022) French CrowS-Pairs: Extending a challenge dataset for measuring social bias in masked language models to a language other than English 1,679 sentence pairs Bias CamemBERTbase, FlauBERTbase, FrALBERT No (French) No
 Cheng et al. (2022b) HiTab: A Hierarchical Table Dataset for Question Answering and Natural Language Generation 3,597 tables QA, NLG Neural Symbolic Machine, TaPas, MML, REINFORCE, MAPO No No
 Cheng et al. (2022a) IAM: A Comprehensive and Large-Scale Dataset for Integrated Argument Mining Tasks 69,666 sentences Argument Mining RoBERTa-base No No
 Cao et al. (2022) KQA Pro: A Dataset with Explicit Compositional Programs for Complex Question Answering over Knowledge Base 117,970 sentences QA KVMemNet, SRN, EmbedKGQA, RGCN, BART No No
 Chalkidis et al. (2022) LexGLUE: A Benchmark Dataset for Legal Language Understanding in English Combination of existing datasets Legal NLU TFIDF-SVM, BERT, RoBERTa, DeBERTa, Longformer, BigBird, Legal-BERT, CaseLaw-BERT No No
 Liang et al. (2022) MSCTD: A Multimodal Sentiment Chat Translation Dataset 142,871 English-Chinese utterance pairs Multimodal chat translation; dialogue sentiment analysis many Yes (English, Chinese) Yes
 Kamezawa et al. (2022) RNSum: A Large-Scale Dataset for Automatic Release Note Generation via Commit Logs Summarization 82,000 English release notes Release Note Generation Glyph, Clustering, BERT, CodeBERT, BART No No
 Chen et al. (2022) SummScreen: A Dataset for Abstractive Screenplay Summarization 26,900 instances Abstractive Summarization Longformer, standard Transformer, BART-large, BM25, ROUGE No No
 Wang et al. (2022c) Textomics: A Dataset for Genomics Data Summary Generation 22,273 pairs of genomics data Genomics Data Summary Generation kNN-Vec2Text, SPECTER, SciBERT, BioBERT, SentBERT, BERT, SentBERT No No
 Hartvigsen et al. (2022) TOXIGEN: A Large-Scale Machine-Generated Dataset for Adversarial and Implicit Hate Speech Detection 274k toxic and benign statements Hate Speech Detection HateBert, ToxDectRoBERTa No No
 Khan et al. (2022) WatClaimCheck: A new Dataset for Claim Entailment and Inference 33,721 claims Automated fact checking TF-IDF, BERT, Bi-LSTM and Bi-GRU, hierarchical attention networks, RoBERTa-base No No
 Wang et al. (2022f) WikiDiverse: A Multimodal Entity Linking Dataset with Diversified Contextual Topics and Entity Types 8K captions Multimodal Entity Linking, NER BM25, WikiVec, BLINK, CLIP, BERT, UNITER, UNITER*, LXMERT No Yes
Table 10: This table shows details for the name of the paper, the size of the dataset, the NLP topic of interest, baselines and multilingualism or multimodality status of the dataset (Part I).
Paper/ Dataset Size (#sentences) NLP Topic Baselines Multi-lingual? Multi-modal?
Track: ACL 2022 continued
 Wang et al. (2022f) WikiDiverse: A Multimodal Entity Linking Dataset with Diversified Contextual Topics and Entity Types 8K captions Multimodal Entity Linking, NER BM25, WikiVec, BLINK, CLIP, BERT, UNITER, UNITER*, LXMERT No Yes
 Filighera et al. (2022) Your Answer is Incorrect… Would you like to know why? Introducing a Bilingual Short Answer Feedback Dataset 4,519 submissions Explainable feedback systems T5-base, mT5-base models Yes (English, German) No
 Bhartiya et al. (2022) DiS-ReX: A Multilingual Dataset for Distantly Supervised Relation Extraction 1.8 million sentences Multilingual relation extraction mBERT Yes (English, French, Spanish, German) No
 Papadopoulos Korfiatis et al. (2022) PriMock57: A Dataset Of Primary Care Mock Consultations 56 mock consultations Automatic Speech Recognition ASR Benchmarks (Kaldi, NeMo QuartzNet, Conformer, Google Cloud Speech-to-text, Amazon Transcribe Medical, Azure Speech-to-text) Consultation Note Generation Benchmarks (BART-CNN, BERT-ext, Random, BART-finet) No No
 Tavella et al. (2022) WLASL-LEX: a Dataset for Recognising Phonological Properties in American Sign Language Sign language recognition MLP, LSTM, GRU, I3D 3D Convolutional Neural Network, Spatio-Temporal Graph Convolutional Network No No
Track: Findings of ACL 2022
 Jiang et al. (2022) Chinese Synesthesia Detection: New Dataset and Models 7,217 sentences synesthesia detection BiLSTM+CRF, BERT+CRF, SR-BiLSTM, PF-BERT No (Chinese) No
 Qi et al. (2022) DuReader_vis: A Chinese Dataset for Open-domain Document Visual Question Answering 15K QA pairs; 158,000 document images Visual Question Answering DocVRE->BM25, DocVQA->RobertaXLM-base, BERT-base-Chinese No (Chinese) Yes
 Ma et al. (2022) EnCBP: A New Benchmark Dataset for Finer-Grained Cultural Background Prediction in English 18K documents cultural background prediction many No No
 Nakamura et al. (2022) HybriDialogue: An Information-Seeking Dialogue Dataset Grounded on Tabular and Textual Data 4,844 dialogues Dialogue Okapi BM25 retriever, SentenceBERT, TaPas, DialoGPT No No
 Yao et al. (2022b) LEVEN: A Large-Scale Chinese Legal Event Detection Dataset 63,616 sentences Legal Event Detection BiLSTM, BiLSTM+CRF, BERT, BERT+CRF, DMCNN, DMBERT No (Chinese) No
 Maronikolakis et al. (2022) Listening to Affected Communities to Define Extreme Speech: Dataset and Experiments 20,297 social media passages hate speech detection mBERT, XLM-R, langBERT Yes (Brazilian Portuguese, German, Hindi, Swahili, English) No
Table 11: This table shows details for the name of the paper, the size of the dataset, the NLP topic of interest, baselines and multilingualism or multimodality status of the dataset (Part II).
Paper/ Dataset Size (#sentences) NLP Topic Baselines Multi-lingual? Multi-modal?
Track: Findings of ACL 2022 continued
 Shen et al. (2022) MReD: A Meta-Review Dataset for Structure-Controllable Text Generation 45k meta-review sentences Controllable Text Generation Extractive Baselines (LexRank, TextRank, MMR, n LSTM-CRF, bart-large-cnn) No No
 Sun et al. (2022b) On the Safety of Conversational Models: Taxonomy, Dataset, and Benchmark 11,492 utterances Dialogue safety DialoGPT, Blenderbot, Plato-2, PerspectiveAPI, Detoxify, Build it, Break it Fix it, BotAdversarial Dialogue Safety Classifier No No
 Xu et al. (2022a) XFUND: A Benchmark Dataset for Multilingual Visually Rich Form Understanding 1,393 forms multilingual document understanding, Semantic Entity Recognition, Relation Extraction LayoutXLM, XLM-RoBERTa, InfoXLM 7 languages (Chinese, Japanese, Spanish, French, Italian, German, Portuguese) Yes
Track: EMNLP 2022
 Thapliyal et al. (2022) Crossmodal-3600: A Massively Multilingual Multimodal Evaluation Dataset 3600 images; 261,375 captions multilingual image captioning ViT, mT5, CIDEr 36 languages Yes
 Eirew et al. (2022) Cross-document Event Coreference Search: Task, Dataset and Modeling 2,772,968 passages Coreference Resolution; event coreference Deep Passage Retrieval (DPR) No No
 Wang et al. (2022e) MAVEN-ERE: A Unified Large-scale Dataset for Event Coreference, Temporal, Causal, and Subevent Relation Extraction 103,193 event coreference chains, 1,216,217 temporal relations, 57,992 causal relations, and 15,841 subevent relations Event Coreference, Temporal, Causal, and Subevent Relation Extraction RoBERTa-BASE No No
 Reid et al. (2022) M2D2: A Massively Multi-Domain Language Modeling Dataset 8.5B tokens Multi-Domain Language Modeling 112M GPT2 No No
 Wang et al. (2022a) SQuALITY: Building a Long-Document Summarization Dataset the Hard Way 100 stories, 625 examples Summarization BART-large, Dense Passage Retriever, PEGASUS, Longformer Encoder-Decoder (LED-base), Human Evaluation No No
 Laban et al. (2022) Near-Negative Distinction: Giving a Second Life to Human Evaluation Datasets N/A NLG evaluation GLUE, GLGE, GEM, BARTScore, No No
 Yao et al. (2022a) D4: a Chinese Dialogue Dataset for Depression-Diagnosis-Oriented Chat 1339 dialogues Dialogue (Depression-Diagnosis-Oriented Chat) BART, CPT, BERT No (Chinese) No
 Liu et al. (2022d) Measuring Context-Word Biases in Lexical Semantic Datasets N/A Model bias BERT No No
 Currey et al. (2022) MT-GenEval: A Counterfactual and Contextual Dataset for Evaluating Gender Accuracy in Machine Translation 300 sentences Gender accuracy in translation Gender Evaluation Benchmarks— WinoMT, MuST-SHE, GeBioCorpus, SimpleGEN. Model— Transformer-base Yes (Arabic, French, German, Hindi, Italian, Portuguese, Russian, Spanish.) No
Table 12: This table shows details for the name of the paper, the size of the dataset, the NLP topic of interest, baselines and multilingualism or multimodality status of the dataset (Part III).
Paper/ Dataset Size (#sentences) NLP Topic Baselines Multi-lingual? Multi-modal?
Track: EMNLP 2022 continued
 Kim et al. (2022b) BotsTalk: Machine-sourced Framework for Automatic Curation of Large-scale Multi-skill Dialogue Datasets 300K conversations Dialogues, open-domain chatbots 256-M parameter poly-encoder, BART, Human Evaluation No No
 Kumar et al. (2022) IndicNLG Benchmark: Multilingual Datasets for Diverse NLG Tasks in Indic Languages 8M examples across 5 tasks (biography generation using Wikipedia infoboxes, news headline generation, sentence summarization, paraphrase generation and, question generation) and 11 languages Multilingual NLG IndicBART, mBART, YANMTT toolkit Yes. 11 languages from o two language families i.e., Indo-Aryan and Dravidian as, bn, gu,hi, kn, ml, mr, or, pa, ta, te No
 Sun et al. (2022d) PHEE: A Dataset for Pharmacovigilance Event Extraction from Text 5000 events drug safety ACE for NER, EEQA for Extractive QA, BioBERT, SciFive for Generative QA, T5 No No
 Guo et al. (2022) Questioning the Validity of Summarization Datasets and Improving Their Factual Consistency 306,522 samples abstractive summarization BART-base No No
 Kim et al. (2022a) Simple Questions Generate Named Entity Recognition Datasets N/A Framework to generate NER datasets RoBERTa, BOND, TALLOR, GeNER No No
 Deng et al. (2022) Title2Event: Benchmarking Open Event Extraction with a Large-scale Chinese Title Dataset 42,000 news titles in 34 topics Event extraction BERT No (Chinese) No
 Liu et al. (2022b) FigMemes: A Dataset for Figurative Language Identification in Politically-Opinionated Memes 5,141 labels figurative language classification in politically-opinionated memes SentenceBERT, Text only (BERT, DeBERTa) Image only (ConvNeXt, CLIP-CNN), Multimodal (CLIP, VinVL, BERT+CLIP, CLIP-MM-OOD) No Yes
 Wang et al. (2022d) ParaTag: A Dataset of Paraphrase Tagging for Fine-Grained Labels, NLG Evaluation, and Data Augmentation 30k sentence pairs Paraphrase identification and generation (1) Paraphrasing detection — BERT, RoBERTa, DeBERTa-v3 (2) Paraphrasing generation — BART No No
 Aumiller et al. (2022) EUR-Lex-Sum: A Multi- and Cross-lingual Dataset for Long-form Summarization in the Legal Domain 1,500 document/summary pairs per language Extractive Summarization Zero-shot Extractive Baselines—LexRank AND Cross-lingual Baselines—Longformer Encoder Decoder (LED) AND translation—OPUS-MT Yes (24 official European languages) No
 Wang et al. (2022b) ClidSum: A Benchmark Dataset for Cross-Lingual Dialogue Summarization 67k+ dialogues AND 112k+ summaries Dialogue Summarization mBART-50 Yes (English, German and Chinese) No
 Zhu et al. (2022) DuQM: A Chinese Dataset of Linguistically Perturbed Natural Questions for Evaluating the Robustness of Question Matching Models 10,121 examples Chinese Question Matching BERT-b, ERNIE-b, RoBERTa-b, MacBERT-b No (Chinese) No
 Bonaldi et al. (2022) Human-Machine Collaboration Approaches to Build a Dialogue Dataset for Hate Speech Countering 3059 dialogues Hate Speech in Dialogue DialoGPT, T5-2m, T5-1m English No
Table 13: This table shows details for the name of the paper, the size of the dataset, the NLP topic of interest, baselines and multilingualism or multimodality status of the dataset (Part IV).
Paper/ Dataset Size (#sentences) NLP Topic Baselines Multi-lingual? Multi-modal?
Track: EMNLP 2022 continued
 Leong et al. (2022) Bloom Library: Multimodal Datasets in 300+ Languages for a Variety of Downstream Tasks Language modeling, Image captioning, Visual storytelling, and Speech synthesis/recognition Language Modeling—DistilBERT, RoBERTa; Image Captioning— InceptionV3; Speech Recognition— Wav2Vec2 XLS-R Yes. 363 languages across 32 language families Yes
 Ravichander et al. (2022) CONDAQA: A Contrastive Reading Comprehension Dataset for Reasoning about Negation 14,182 question-answer pairs Reading Comprehension, Reasoning Full fine-tuning (BERT, RoBERTa, DeBERTa, T5 variant namely UnifiedQA-v2-3b) AND Few-shot setting (UnifiedQA-v2-Base, Large, 3B , InstructGPT-orig ie. text-davinci-002) and Human Evaluation No No
 Smith et al. (2022) “I’m sorry to hear that”: Finding New Biases in Language Models with a Holistic Descriptor Dataset 459,758 sentences Bias GPT2-large, RoBERTa, DialoGPT, BlenderBot 2.0 No No
 Pouran Ben Veyseh et al. (2022) MEE: A Novel Multilingual Event Extraction Dataset 50K event mentions Event Extraction Pipeline approach — OneIE, FourIE; Joint approach— mBERT, XLM-RoBERTa Yes. 8 languages (English, Spanish, Portuguese, Polish, Turkish, Hindi, Korean, and Japanese) No
 Mekala et al. (2022) Leveraging QA Datasets to Improve Generative Data Augmentation N/A Generate synthetic data (Data Augmentation) BERT-base, BackTranslation, PEGASUS, EDA, LAMBADA No No
 Chia et al. (2022) A Dataset for Hyper-Relational Extraction and a Cube-Filling Approach 44,372 unique facts Relational Extraction BERT, Feed-forward network. Pipeline Baseline—UniRE, BERT-Tagger, DistilBERT. Generative Baseline— BART No No
 Zhao et al. (2022b) A Fine-grained Chinese Software Privacy Policy Dataset for Sequence Labeling and Regulation Compliant Identification 11K sentences, 52K fine-grained annotations Sequence Labeling Hidden Markov Model (HMM), Conditional Random Field (CRF), BiLSTM, BiLSTM-CRF, BERT-BiLSTM-CRF No (chinese) No
 Jeong et al. (2022) KOLD: Korean Offensive Language Dataset 40,429 comments Offensive Language detection Korean BERT and RoBERTa models No (Korean) No
 Mukherjee et al. (2022) ECTSum: A New Benchmark Dataset For Bullet Point Summarization of Long Earnings Call Transcripts 2,425 document-summary pairs Summarization FinBERT, T5, LexRank, PacSum , SummaRuNNer, BertSumEXT, MatchSum, BART, Pegasus, BigBird, Longformer Encoder Decoder No No
 Varshney et al. (2022) CDialog: A Multi-turn Covid-19 Conversation Dataset for Entity-Aware Dialog Generation 1,012 dialogs and 7,982 utterances Entity-Aware Dialog Generation BioBERT-BASE, GPT-2, DialogGPT-finetune, BERT, BART, Human Evaluation. No No
 Pujari et al. (2022) Three Real-World Datasets and Neural Computational Models for Classification Tasks in Patent Landscaping 9,465 instances (Injection-Valve) Patent Classification SciBERT, Transformer-based Multi-task Model No No
Table 14: This table shows details for the name of the paper, the size of the dataset, the NLP topic of interest, baselines and multilingualism or multimodality status of the dataset (Part V).
Paper/ Dataset Size (#sentences) NLP Topic Baselines Multi-lingual? Multi-modal?
Track: EMNLP 2022 continued
 Ye et al. (2022a) ZeroGen: Efficient Zero-shot Learning via Dataset Generation N/A Zero-shot Learning GPT2, GPT2-large, GPT2-XL, LSTM, DistilBERT No No
 Zhao et al. (2022c) JDDC 2.1: A Multimodal Chinese Dialogue Dataset with Joint Tasks of Query Rewriting, Response Generation, Discourse Parsing, and Summarization 246K dialogue sessions, 3M utterances, and 507K images multimodal dialogue response generation task, query rewriting task, dialogue discourse parsing task, dialogue summarization BERT, k-NN, ResNet18, GPT-2, TransResNet No (Chinese) Yes
 Eklund and Forsman (2022) Topic Modeling by Clustering Language Model Embeddings: Human Validation on an Industry Dataset N/A Topic Modeling BERT, UMAP, HDBSCAN No No
Track: Findings of EMNLP 2022
 Qu et al. (2022) Commonsense Knowledge Salience Evaluation with a Benchmark Dataset in E-commerce 30K entities and relations salience of commonsense knowledge (CSK) BERTSAGE, KG-BERT, StAR, GenKGC, PKGC, No No
 Zhao et al. (2022a) NarraSum: A Large-Scale Dataset for Abstractive Narrative Summarization 122K narratives Abstractive Summarization RANDOM, LEAD, TEXTRANK, LEXRANK, HSG, PRESUMM; Extractive summarization — BERT, ROBERTA, LONGFORMER; Abstractive summarization — BART, T5, PEGASUS, LED No No
 Zhang et al. (2022) MovieUN: A Dataset for Movie Understanding and Narrating TNG task— 3,253 long video clips and MCN task— 33,060 video clips Movie Understanding and Narrating Vanilla Transformer, OVP, CLIP, MIL-NCE, e Arcface model No (Chinese) Yes
 Liu et al. (2022e) Long Text and Multi-Table Summarization: Dataset and Method 21,125 annual reports from 3,794 companies long text and multi-table summarization LexRank, TextRank, BART, PEGASUS, T5, BigBird-PEGASUS, Longformer Encoder Decoder (LED) No No
 Sun et al. (2022c) Visual Named Entity Linking: A New Dataset and A Baseline 48K images, 120K named entities Visual Entity Linking ResNet, CLIP, No Yes
 Ekram et al. (2022) BanglaRQA: A Benchmark Dataset for Under-resourced Bangla Language Reading Comprehension-based Question Answering with Diverse Question-Answer Types 14,889 question-answer pairs QA BanglaT5, mT5, BanglaBERT, mBERT No (Bangla) No
 Zhou et al. (2022) Towards Identifying Social Bias in Dialog Systems: Framework, Dataset, and Benchmark 28K context-response pairs Dialog systems Bert-Base-Chinese, VANILLA, MIXTURE-OF EXPERTS, CDIAL-GPT, EVA, CPM No (Chinese) No
 Bassignana and Plank (2022) CrossRE: A Cross-Domain Dataset for Relation Extraction 5,265 sentences and 18,608 relations Relation Extraction BERT No No
 Ye et al. (2022b) ProGen: Progressive Zero-shot Dataset Generation via In-context Feedback Dataset Generation LSTM, DistilBERT ?? ??
 Otmakhova et al. (2022) M3: Multi-level dataset for Multi-document summarisation of Medical studies 6365 document Multi-document summarization BART, BioBART, Pegasus, BigBird Pegasus, Primera No No
Table 15: This table shows details for the name of the paper, the size of the dataset, the NLP topic of interest, baselines and multilingualism or multimodality status of the dataset (Part VI).
Paper/ Dataset Size (#sentences) NLP Topic Baselines Multi-lingual? Multi-modal?
Track: Findings of EMNLP 2022 continued
 Fekih et al. (2022) HumSet: Dataset of Multilingual Information Extraction and Classification for Humanitarian Crises Response 148,621 entries and 62 different tags Multilingual Information Extraction Entry extraction— LEAD4, XtremeDistil, XLM-R-Base and Entry classification— FastText, XtremeDistil, XLM-R-Base Yes (English, French, Spanish) No
 Bastan et al. (2022) BioNLI: Generating a Biomedical NLI Dataset Using Lexico-semantic Constraints for Adversarial Examples 21997 instances NLI PubMedBERT, BioLinkBERT, No No
 Armstrong et al. (2022) JamPatoisNLI: A Jamaican Patois Natural Language Inference Dataset 650 examples Patois NLI English BERT, multilingual BERT, English RoBERTa, XLM-RoBERTa No (Patois) No
 Lai et al. (2022) Multilingual SubEvent Relation Extraction: A Novel Dataset and Structure Induction Method 3,591 documents Relation Extraction RoBERTa-base, mBERT, XLMR, StructLR, TacoLM, Joint, EventSeg, SCS Yes. 5 languages (English, Danish, Spanish, Turkish, and Urdu) No
 Liu et al. (2022c) Augmenting Multi-Turn Text-to-SQL Datasets with Self-Play Text-to-SQL PICARD/ T5-Base, T5-Large No No
 Srinivasan and Choi (2022) TyDiP: A Dataset for Politeness Classification in Nine Typologically Diverse Languages 4.5K examples Politeness Classification English-RoBERTa, XLM-RoBERTa, GPT3 Davinci-002 Yes. 9 languages (Hindi, Korean, Spanish, Tamil, French, Vietnamese, Russian, Afrikaans, and Hungarian.) No
 Yang et al. (2022b) PcMSP: A Dataset for Scientific Action Graphs Extraction from Polycrystalline Materials Synthesis Procedure Text 305 documents, 2468 sentences, 14592 entities, 13968 relations Action Graphs Extraction BERT-base, SciBERT, MatBERT No No
 Liu et al. (2022f) Code Generation From Flowcharts with Texts: A Benchmark Dataset and An Approach 320 flowcharts source code generation LSTM, Transformer No No
 Maheshwari et al. (2022) A Benchmark and Dataset for Post-OCR text correction in Sanskrit 218,000 sentences, Post-OCR text correction CopyNet, mBART, mT5, ByT5, IndicBART No (Sanskrit) No
 Kryscinski et al. (2022) BOOKSUM: A Collection of Datasets for Long-form Narrative Summarization 405 documents, Summarization Lead-3, CNN-LSTM Extractor, BertExt, MatchSum, BART, T5, PEGASUS No No
 Liu et al. (2022a) WANLI: Worker and AI Collaboration for Natural Language Inference Dataset Creation 107,885 examples NLI RoBERTa-large, GPT-3 No No
 Yang et al. (2022a) APEACH: Attacking Pejorative Expressions with Analysis on Crowd-Generated Hate Speech Evaluation Datasets hate speech detection KoBERT, DistilKoBERT, KoELECTRA, KcBERT, SoongsilBERT No (Korean) No
 Thapa et al. (2022) A Multi-Modal Dataset for Hate Speech Detection on Social Media: Case-study of Russia-Ukraine Conflict 5,680 text-image pairs Hate Speech Detection ResNet+BERT, VisualBERT, CLIP; Text unimodal—LSTM, BERT, RoBERTa, and Visual unimodal—DenseNet, ResNet, VGG-19 No Yes
 Sticha and Brenner (2022) Hybrid Knowledge Engineering Leveraging a Robust ML Framework to Produce an Assassination Dataset 7,457 assassination events Socio-policial event Extraction SVM, BERT, Longformer, No No
 Wiriyathammabhum (2022) ClassBases at the CASE-2022 Multilingual Protest Event Detection Task: Multilingual Protest News Detection and Automatically Replicating Manually Created Event Datasets COVID-related protest event datasets from the New York Times news corpus event detection document classification—XLM-RoBERTa-base, mLUKE-base and XLM-RoBERTa-large AND sentence classification— mLUKE-base AND token classification—XLM-RoBERTa-base Yes (English, Spanish and Portuguese) No
Table 16: This table shows details for the name of the paper, the size of the dataset, the NLP topic of interest, baselines and multilingualism or multimodality status of the dataset (Part VII).
Appendix C BLOOM Dataset
Table 17 shows all the languages represented in the BLOOM dataset.
The 363 Languages in BLOOM Library
Ghotuo[aaa]; Ayta, Ambala[abc]; Dangme[ada]; Adangbe[adq]; Akeu[aeu]; Afrikaans[afr]; Aghem[agq]; Esimbi[ags]; Akha[ahk]; Arosi[aia]; Amri Karbi[ajz]; Akan[aka]; Yanesha’[ame]; Amharic[amh]; Alamblak[amp]; Amuzgo, Guerrero[amu]; Obolo[ann]; Athpariya[aph]; Awadhi[awa]; Awa[awb]; Nahuatl, Western Durango[azn]; Awing[azo]; Tuki[bag]; Bamanankan[bam]; Bambili-Bambui[baw]; Bamun[bax]; Babanki[bbk]; Balochi, Southern[bcc]; Bamenyam[bce]; Iceve-Maci[bec]; Benabena[bef]; Bengali[ben]; Bafut[bfd]; Mmen[bfm]; Bunak[bfn]; Bangandu[bgf]; Bhojpuri[bho]; Buwal[bhs]; Bislama[bis]; Banjar[bjn]; Binumarien[bjr]; Baka[bkc]; Bakoko[bkh]; Kom[bkm]; Baikeno[bkx]; Aweer[bob]; Tibetan, Central[bod]; Bozo, Tieyaxo[boz]; Wumboko[bqm]; Braj Bhasha[bra]; Lave[brb]; Mokpwe[bri]; Bru, Western[brv]; Akoose[bss]; Ntcham[bud]; Terei[buo]; Bafaw-Balong[bwt]; Bunu, Bu-Nao[bwx]; Tairaha[bxa]; Bukusu[bxk]; Batak[bya]; Bozo, Jenaama[bze]; Bisu[bzi]; Kaqchikel[cak]; Kakataibo-Kashibo[cbr]; Cebuano[ceb]; Kagayanen[cgc]; Chontal, Highland Oaxaca[chd]; Dene[chp]; Cimbrian[cim]; Kurdish, Central[ckb]; Chontal, Lowland Oaxaca[clo]; Chinese, Mandarin[cmn]; Chinese, Mandarin[cmn]; Mnong, Central[cmo]; Cree, Swampy[csw]; Gichuka[cuh]; Cuvok[cuv]; Dagbani[dag]; Fataluku[ddg]; Dedua[ded]; German, Standard[deu]; Chidigo[dig]; Zarma[dje]; Kinabatangan, Upper[dmg]; Dani, Western[dnw]; Kadazan Dusun[dtp]; Lotud[dtr]; Dotyali[dty]; Chiduruma[dug]; Elip[ekm]; Markweeta[enb]; En[enc]; English[eng]; Ewondo[ewo]; Filipino[fil]; Fali[fli]; Fon[fon]; French[fra]; Fulfulde, Adamawa[fub]; Fulfulde, Western Niger[fuh]; Galolen[gal]; Gadaba, Bodo[gbj]; Gavar[gou]; German, Swiss[gsw]; Wayuu[guc]; Gujarati[guj]; Ekegusii[guz]; Gawri[gwc]; Hakö[hao]; Haitian Creole[hat]; Hausa[hau]; Nya Huba[hbb]; Kamwe[hig]; Hiligaynon[hil]; Hindi[hin]; Halia[hla]; Mina[hna]; Hre[hre]; Haroi[hro]; Idaté[idt]; Ilocano[ilo]; Indonesian[ind]; Inoke-Yate[ino]; Isu[isu]; Italian[ita]; Ngomba[jgo]; Mixtec, Western Juxtlahuaca[jmx]; Japanese[jpn]; Jarai[jra]; Kalanguya[kak]; Kamba[kam]; Kannada[kan]; Kamano[kbq]; Ap Ma[kbx]; Kanuri, Manga[kby]; Kanuri, Manga[kby]; Q’eqchi’[kek]; Kenyang[ken]; Lü[khb]; Khmer[khm]; Gikuyu[kik]; Kinyarwanda[kin]; Kyrgyz[kir]; Q’anjob’al[kjb]; Kâte[kmg]; Kurdish, Northern[kmr]; Kamasau[kms]; Kanite[kmu]; Korean[kor]; Kimaragang[kqr]; Krung[krr]; Karen, S’gaw[ksw]; Lahta[kvt]; Kwaio[kwd]; Kwakum[kwu]; Khirwar[kwx]; Koli, Wadiyari[kxp]; Kenga[kyq]; Lango[laj]; Laru[lan]; Lao[lao]; Lohorung[lbr]; Lefa[lfa]; Lugbara[lgg]; Lengo[lgr]; Lhomi[lhm]; Lahu[lhu]; Lukabaras[lkb]; Lole[llg]; Limbum[lmp]; Lamnso’[lns]; Narim[loh]; Lacid[lsi]; Lutachoni[lts]; Ganda[lug]; Lawa, Eastern[lwl]; Maithili[mai]; Malayalam[mal]; Mam[mam]; Marathi[mar]; Mandar[mdr]; Matal[mfh]; Mefele[mfj]; Mpumpong[mgg]; Mambae[mgm]; Meta’[mgo]; Malila[mgq]; Lhao Vo[mhx]; Mixtec, Ayutla[miy]; Makasae[mkz]; Manambu[mle]; Kiwilwana[mlk]; Moloko[mlw]; Mmaala[mmu]; Naba[mne]; Mundani[mnf]; Mon[mnw]; Barí[mot]; Mamasa[mqj]; Cheke Holo[mrn]; Mandaya[mry]; Masbatenyo[msb]; Muthuvan[muv]; Marwari[mve]; Mada[mxu]; Burmese[mya]; Sénoufo, Mamara[myk]; Masaaba[myx]; Mumuye[mzm]; Naasioi[nas]; Sibe[nco]; Newar[new]; Ngemba[nge]; Ngwo[ngn]; Nahuatl, Isthmus-Mecayapan[nhx]; Njyem[njy]; Ngombale[nla]; Dutch[nld]; Nahuatl, Orizaba[nlv]; Thai, Northern[nod]; Nepali[npi]; Naskapi[nsk]; Nehan[nsn]; Sotho, Northern[nso]; Naga, Tangshang[nst]; Nyole[nuj]; Ngwe[nwe]; Tanna, Southwest[nwi]; Nauete[nxa]; Nuaulu, South[nxl]; Chichewa[nya]; Nyoro[nyo]; Nyungwe[nyu]; Mbembe, Tigon[nza]; Oadki[odk]; Oji-Cree[ojs]; Okiek[oki]; Tairora, South[omw]; Odia[ory]; Koonzime[ozm]; Pagibete[pae]; Pangasinan[pag]; Punjabi, Eastern[pan]; Pashto, Southern[pbt]; Palaung, Ruching[pce]; Paniya[pcg]; Kayan[pdu]; Indonesian, Peranakan[pea]; Persian, Iranian[pes]; Petats[pex]; Pijin[pis]; Kipfokomo[pkb]; Pamona[pmf]; Pana[pnz]; Portuguese[por]; Gapapaiwa[pwg]; Quechua, Huallaga[qub]; K’iche’[quc]; Quechua, Lambayeque[quf]; Quechua, Cusco[quz]; Quechua, Eastern Apurímac[qve]; Quechua, Huamalíes-Dos de Mayo Huánuco[qvh]; Quechua, Margos-Yarowilca-Lauricocha[qvm]; Quichua, Napo[qvo]; Quechua, Panao[qxh]; Rendille[rel]; Ranglong[rnl]; Romanian[ron]; Rotokas[roo]; Rusyn[rue]; Roviana[rug]; Russian[rus]; Sanskrit[san]; Samburu[saq]; Santhali[sat]; Sos Kundi[sdk]; Semai[sea]; Surigaonon[sgd]; Shan[shn]; Sama, Central[sml]; Soninke[snk]; Sangil[snl]; Somali[som]; Sotho, Southern[sot]; Swo[sox]; Spanish[spa]; Saposa[sps]; Waata[ssn]; Arammba[stk]; Swahili[swh]; Swahili[swh]; Suba[sxb]; Syuba[syw]; Tamang, Eastern[taj]; Tamil[tam]; Tiang[tbj]; Panchpargania[tdb]; Emberá-Tadó[tdc]; Tamang, Western[tdg]; Tetun Dili[tdt]; Teso[teo]; Tetun[tet]; Tajik[tgk]; Tagalog[tgl]; Thai[tha]; Tharu, Madhya Ksetriya[the]; Kitharaka[thk]; Tharu, Dangaura[thl]; Tha[thy]; Teop[tio]; Tukudede[tkd]; Lenakel[tnl]; Tanna, North[tnn]; Whitesands[tnp]; Tontemboan[tnt]; Toma[tod]; Tombulu[tom]; Tok Pisin[tpi]; Me’phaa, Tlacoapa[tpl]; Tampuan[tpu]; Tsamai[tsb]; Setswana[tsn]; Tsonga[tso]; Turkana[tuv]; Turka[tuz]; Taveta[tvs]; Tz’utujil[tzj]; Muduga[udg]; Mundari[unr]; Urdu[urd]; Uzbek, Northern[uzn]; Venda[ven]; Vietnamese[vie]; Vili[vif]; Waray-Waray[war]; Wa, Vo[wbm]; Wagdi[wbr]; Wambon[wms]; Comorian, Ndzwani[wni]; Wanukaka[wnk]; Watakataui[wtk]; Xhosa[xho]; Kagoro[xkg]; Mbudum[xmd]; Mengaka[xmg]; Malay, Manado[xmm]; Soga[xog]; Mixtec, Yoloxóchitl[xty]; Nugunu[yas]; Yangben[yav]; Yemba[ybb]; Yakkha[ybh]; Yamphu[ybi]; Yiddish, Eastern[ydd]; Yiddish, Eastern[ydd]; Ravula[yea]; Riang Lai[yin]; Yamap[ymp]; Zapotec, Mitla[zaw]; Malay[zlm]; Tokano[zuh]; Zulu[zul]
Table 17: The full list of languages in the BLOOM (Leong et al., 2022) dataset.